# Sylva: Tailoring Personalized Adversarial Defense in Pre-trained Models via Collaborative Fine-tuning

### v1

@@ 4.3. Performance of Different Heterogeneous Distributions (RQ2)
Table 3. Detailed specifications of GPUs in real edge devices
@@
Figure 8. Commonly used GPUs in autonomous driving scenarios
Table 4. Efficiency comparison of Sylva on different edge devices
[…]
ViT-T FAT 1.04 0.35 0.45 0.47 0.49 0.99 0.7 1.3 1.5 1.0 1.2
Per-LoRA 0.89 0.31 0.32 0.33 0.32 0.82 0.9 0.7 1.1 1.1 1.0
Sylva 0.89 0.30 0.32 0.33 0.32 0.82 0.8 0.6 1.2 0.9 0.9
ViT-B FAT 4.86 0.40 0.47 0.49 0.53 1.13 15.0 13.5 13.8 14.1 13.8
Per-LoRA 3.14 0.34 0.39 0.42 0.45 0.92 1.3 1.4 1.4 1.2 1.5
Sylva 3.14 0.33 0.37 0.41 0.45 0.91 0.8 0.9 0.8 0.7 0.8
ViT-L FAT 11.43 0.75 0.93 1.78 1.87 2.41 49.9 47.6 49.1 48.7 48.2
Per-LoRA 6.63 0.57 0.88 1.29 1.39 1.71 3.1 2.9 3.2 3.4 3.0
Sylva 6.63 0.55 0.87 1.28 1.39 1.70 1.1 1.0 1.2 1.2 1.1
(a) CIFAR-10
(b) STL-10
@@
Figure 9. Impact of different loss modules on adversarial robustness and benign accuracy on CIFAR-10 and STL-10
4.4. Comparison of Adversarial Training Efficiency (RQ3)
Given that Sylva is designed for deployment in distributed systems with multiple edge devices, evaluating its efficiency in real-world environments is essential. To simulate a real distributed system, we employ three widely used GPUs—NVIDIA GeForce RTX 4090, 3090, and 2080-Ti—on edge devices, while leveraging a high-speed subnet server as the cloud. To align with real-world scenarios, we also explore the deployment of Sylva in autonomous driving, including tasks like road sign recognition, as described in our threat model. We test Sylva on both the autonomous driving simulation device, specifically the RTX 3060 in the Apollo D-KIT Advanced (Baidu, 2022), and the widely used Jetson AGX Orin in real vehicles (NVIDIA, 2021). The specifications of the GPUs used are shown in Table 3. We present key GPU parameters, including memory size, clock frequency, bandwidth, and the 16-bit floating-point (FP16) tera floating-point operations per second (TFLOPS), which are crucial for evaluating computational performance, with TFLOPS serving as a key indicator of computational power. Based on these parameters, theoretical analysis indicates that the consumer-grade GPUs—RTX 4090, 3090, and 2080-Ti—deliver excellent performance. In autonomous driving scenarios, the RTX 3060 performs slightly less effectively, as the Apollo D-KIT Advanced is primarily designed for vehicle task simulation. In contrast, real-world vehicles are typically equipped with 2-4 Jetson AGX Orin units (NVIDIA, 2024a, b), allowing parallel deployment to approach the computational power of the RTX 3090 per vehicle.
In this setup, we assess Sylva’s efficiency by measuring its memory overhead (Mem), per-epoch training time (Time), and model upload communication time (Com) on edge devices. These metrics are compared against FA-TRADES (FAT) to provide a comprehensive evaluation. The results are shown in Table 4. Compared to traditional FAT algorithms, Sylva employs an efficient parameter fine-tuning approach for pre-trained models, offering significant reductions in memory usage, training time, and communication overhead. These benefits become more pronounced as model size increases. Specifically, Sylva reduces memory usage by up to 35.6%, training time by 26.7% on high-performance GPUs like the NVIDIA 4090, and up to 28.1% on resource-constrained devices such as the 2080-Ti. In terms of communication, Sylva achieves a 50$\times$ reduction in communication time, even with larger models and high-speed local networks, making it ideal for real-world distributed systems. Compared to Per-LoRA, Sylva shows no significant difference in memory usage, as both use LoRA for fine-tuning. However, Sylva’s unique personalization approach only requires the aggregation of the backbone’s LoRA parameters, while Per-LoRA also needs to upload the classifier’s parameters. This results in a communication savings of up to 64.5% for Sylva. Additionally, Per-LoRA’s trust weight computation adds complexity, while Sylva’s ball tree aggregation reduces complexity, leading to slightly faster training.
On the other hand, Sylva not only improves performance on high-end GPUs but also reduces training time and communication overhead on real edge devices. Experimental results closely match the computational power estimates derived from our theoretical analysis based on GPU parameters. Notably, for the Jetson AGX Orin, we only use a single unit for the experiments, but real-world vehicles typically use multiple units. Model or data parallelism could potentially double the training speed (Fan et al., 2021). With 32GB of memory, the Orin can run Sylva alongside other autonomous driving tasks. Overall, Sylva outperforms baseline algorithms in real-world scenarios and is viable for practical deployment. In addition, we also test the communication efficiency when training with different numbers of clients, and the results are provided in Appendix D.
Answer to RQ3: Sylva significantly enhances distributed adversarial training by reducing memory usage, training time, and communication overhead, with its advantages growing as model size increases, making it highly efficient for real-world scenarios.
(a) CIFAR-10
(b) STL-10
@@ 4.4. Comparison of Adversarial Training Efficiency (RQ3)
Figure 10. Impact of ball-tree-based aggregation algorithm on CIFAR-10 and STL-10
(a) CIFAR-10
(b) STL-10
@@
Figure 11. Impact of different phases on the trade-off between robustness and accuracy for CIFAR-10 and STL-10
4.5. Ablation Experiments (RQ4)
4.5.1. Impact of Different Modules.
We first evaluate the impact of our loss function, as illustrated in Fig. 9, by testing the total loss function and assessing adversarial robustness and benign accuracy without the $\mathcal{L}_{A}$, $\mathcal{L}_{S}$, and $\mathcal{L}_{R}$ components individually. The results clearly demonstrate that each component plays a positive role in enhancing overall performance. Among these, $\mathcal{L}_{A}$ and $\mathcal{L}_{S}$ prove to be particularly critical, as their absence leads to a substantial decline in both robustness and accuracy, underscoring their importance in the design of the loss function. We next evaluate the impact of the ball tree aggregation algorithm employed in Phase 1 of Sylva by comparing it with the traditional aggregation method used in FedAvg (FA-agg). As shown in Fig. 10, the proposed algorithm effectively mitigates model drift caused by data heterogeneity, resulting in superior performance in both adversarial robustness and benign accuracy. We further evaluate the effectiveness of the two-phase design on model robustness and benign accuracy. Phase 1 (P1) adversarial training generates a highly robust model, establishing a strong foundation for further refinement. In Phase 2 (P2), we compare Sylva’s algorithm with Gen-AF’s Stage 2 method by training the robust models with their respective approaches. As shown in Fig. 11, Phase 2 achieves a substantial improvement in benign accuracy with only a minimal trade-off in adversarial robustness, clearly outperforming the Gen-AF approach. These results highlight the ability of Sylva’s two-phase design to effectively balance adversarial robustness and benign accuracy. The experiments discussed above are conducted on CIFAR-10 and STL-10. Additional results for GTSRB and CIFAR-100 are provided in Appendix E for further validation and analysis.
Table 5. The Impact of Hyperparameters in CIFAR-10

### v3

@@ A-D Data Heterogeneity
TABLE VIII: Detailed specifications of GPUs in real edge devices
@@
Figure 17: Commonly used GPUs in autonomous driving scenarios
(a) ViT-T
(b) ViT-B
(c) ViT-L
Figure 18: Relationship between the number of clients and communication time
table II provides a comprehensive comparison of Lorica against baseline models utilizing the ViT-B architecture across multiple datasets. The results emphasize Lorica’s superior performance in terms of both adversarial robustness and benign accuracy under a wide range of adversarial attack scenarios. Notably, Lorica consistently surpasses the baseline methods, demonstrating its capability to effectively manage diverse attack conditions while maintaining exceptional levels of adversarial reliability.
To further evaluate the adaptability of Lorica, we extend our analysis to models of varying sizes, specifically ViT-T and ViT-L, with detailed results presented in table VI and table VII. The findings highlight that Lorica maintains strong defense capabilities across different model scales, delivering robust performance even as model complexity changes. Lorica outperforms both traditional distributed algorithms and personalized algorithms, consistently demonstrating superior performance in terms of both robustness and adaptability. Notably, while the defense capabilities of baseline methods often degrade significantly with smaller models, such as ViT-T, Lorica demonstrates exceptional adaptability. It consistently achieves an effective balance between adversarial robustness and benign accuracy, underscoring its versatility and resilience under varying conditions. Moreover, Lorica achieves overall better performance than Sylva, both in terms of accuracy and robustness. These findings highlight the remarkable versatility of Lorica, demonstrating its suitability for a broad spectrum of model architectures and adversarial scenarios. Its consistent ability to deliver robust performance across diverse datasets and varying model scales underscores its potential as a dependable solution for adversarial training. This adaptability makes Lorica an effective choice for applications ranging from simple to highly complex environments, further cementing its value in enhancing model resilience.
(a) Label-flipping
(b) MPAF
Figure 19: Defense performance against Byzantine attacks on GTSRB
(a) Label-flipping
(b) MPAF
Figure 20: Defense performance against Byzantine attacks on CIFAR-100
(a) GTSRB
(b) CIFAR-100
@@ Appendix B Additional Results on the Performance of Adversarial Training (RQ1)
Figure 21: Impact of different loss modules on adversarial robustness and benign accuracy on GTSRB and CIFAR-100
In section IV-C, we thoroughly analyze the adversarial defense performance of the ViT-B model under diverse heterogeneity conditions, with Lorica demonstrating remarkable defense capabilities and robustness throughout. To gain deeper insights into how models of varying scales respond to heterogeneity, we extend the evaluation to the ViT-T and ViT-L models, systematically assessing their defense performance under different levels of heterogeneity, as depicted in fig. 15 and fig. 16, respectively.
The results demonstrate that Lorica consistently outperforms its counterparts across various model sizes, effectively maintaining both accuracy and robustness, even in highly challenging heterogeneous environments where data distribution and computational resources vary significantly. Notably, smaller models, such as ViT-T, exhibit more rapid convergence during training, enabling them to achieve relatively stable performance despite the increasing heterogeneity of the data. This finding underscores Lorica’s adaptability in managing the intricate balance between model complexity and environmental variability. Moreover, compared to the other two personalized baseline algorithms, the smaller performance fluctuations observed under heightened heterogeneity indicate that Lorica’s personalized framework effectively mitigates the detrimental impacts of data heterogeneity by tailoring defense strategies to individual clients. In addition, Lorica demonstrates consistently superior performance to Sylva across models of various scales, confirming the effectiveness of its personalization-driven optimization strategy. These findings highlight the exceptional versatility of Lorica, making it an ideal choice for adversarial training across systems characterized by diverse heterogeneity and a broad spectrum of model sizes.
(a) GTSRB
(b) CIFAR-100
@@ Appendix C Additional Results on the Performance Under Different Heterogeneity Levels (RQ2)
Figure 22: Impact of ball-tree-based aggregation algorithm on GTSRB and CIFAR-100
(a) GTSRB
(b) CIFAR-100
@@
Figure 23: Impact of different phases on the trade-off between robustness and accuracy for GTSRB and CIFAR-100
TABLE IX: The Impact of Hyperparameters in STL-10